OpenStudy (anonymous):

Find the intersection point(s) of the following two functions: a(x) = x2 - 3 b(x) = -x2 + 5

OpenStudy (anonymous):

y = x^2 - 3 y = -x^2 + 5 2y = 2 y = 1 1 = x^2 - 3 4 = x^2 x = 2, -2 (2,1) (-2,1) are the intersecting points.

OpenStudy (anonymous):

thank you.